40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 28 17:00:50 2013, Jenne, Update, ASC, Proposed POP path, to be installed this evening CurrentPOPpath.pdfProposedPOPpath.pdf
    Reply  Thu May 30 17:18:58 2013, Jenne, Update, ASC, New POP path 
       Reply  Sat Jun 1 14:14:56 2013, Jenne, Update, ASC, New POP path POX_table_30May2013progress.pdfPOP_QPD_lensSolution.png
          Reply  Fri Jun 14 03:10:40 2013, Jenne, Update, ASC, New POP path - PDs in place, need cabling 
             Reply  Fri Jun 14 17:54:11 2013, Jenne, Update, ASC, New POP path - cabling work 
             Reply  Wed Jun 19 22:02:03 2013, Jenne, Update, ASC, New POP path - ready to try 
Message ID: 8728     Entry time: Wed Jun 19 22:02:03 2013     In reply to: 8707
Author: Jenne 
Type: Update 
Category: ASC 
Subject: New POP path - ready to try 

I put the POPDC cable back to the DC output of the bias tee that is the first thing at the LSC rack that the POP110 PD sees.  So, now we should be back to the old nominal PRCL locking, with the addition of the new QPD.

I'm going to give it a whirl.....

ELOG V3.1.3-